# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v5

@@ V-A7 Constrained learning
While the previously discussed pathways focus on enhancing a robot’s capabilities, a parallel and equally critical frontier is ensuring that these capabilities are exercised safely and reliably. To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [81, 82].
Soft constraints encourage desirable behaviour and penalize undesirable behaviour without forbidding it. They are well suited for preferences or efficiency goals, for example minimizing energy use, limiting peak torques, or promoting smooth motion [81]. Hard constraints are inviolable rules that prevent catastrophic failures. They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [54, 90]. A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [92].
@@ V-B1 Leveraging foundation models for locomotion learning
The recent rise of Foundation Models (FMs), such as Large Language Models (LLMs) and Vision-Language Models (VLMs), presents a transformative opportunity for bipedal locomotion. Their powerful reasoning capabilities are unlocking new approaches that go beyond traditional control methods, primarily by enabling sophisticated HL task planning and by providing novel solutions to shape the learning process itself, particularly in automated reward design.
As HL planners, FMs provide a reasoning engine that can bridge the gap between abstract human goals and LL motor execution. They can interpret complex linguistic commands or visual scenes and decompose them into a sequence of simpler, actionable commands for a LL policy to follow. This has been demonstrated effectively in legged robotics where VLMs process raw sensory data to pass structured commands to motor controllers [93], creating a seamless link between strategic planning and physical action.
Furthermore, FMs create a significant opportunity to overcome one of the most persistent bottlenecks in DRL: reward design. Instead of tedious manual tuning, LLMs can dynamically generate or refine reward functions based on linguistic descriptions of task success. Research has shown that LLMs can translate human feedback into reward adjustments [94] or even autonomously adjust rewards and control strategies to self-optimize for diverse terrains [95], drastically reducing human intervention.
The foremost opportunity lies in the deeper synergy between these roles. The integration of the HL symbolic reasoning of FMs with the LL, real-time control of DRL could create a new class of highly adaptive and flexible robots. As this rapidly evolving field progresses, as reviewed in [96], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
@@ V-B2 Loco-manipulation tasks
While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it can’t physically interact with its environment. The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction. Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [42].
However, realizing this opportunity is a significant challenge. Early studies, such as a ‘box transportation’ framework [97], often rely on inefficient, multi-policy solutions that lack visual perception. Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [98, 99].
These difficulties create significant research opportunities. One such opportunity lies in exploring hierarchical control approaches. By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [56].
Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution. Using techniques like curriculum learning and imitation from human motion capture data [51, 39, 35, 100, 31, 29], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
@@ V-B3 Insights from quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [101, 102], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [103]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [104, 105], and sophisticated acrobatic motions are achieved through IL [106].
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [107] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [62, 108] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs a MPC to generate reference motions, which are then followed by a LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots module [109] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are underway. These include mimicking real animals through motion capture data and IL [110, 111], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [112, 113] or by using the robots’ legs [114]. Notably, the research presented in [113] demonstrates that loco-manipulation tasks can be effectively managed using a single unified end-to-end framework.
Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers and high mechanical complexity can be attributed to the cause of this disparity between quadruped and bipedal robots. Establishing a unified framework could bridge this gap, an essential step, given the integral role of bipedal robots with upper bodies in developing fully functional humanoid systems. Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
@@ V-C Conceptual Models for Unified Frameworks
Motivated by our survey and the current state of the art, we propose two conceptual models, intended as reference designs, toward a unified locomotion framework. They build on end-to-end and hierarchical paradigms and offer complementary routes to scalable, generalizable architectures, rather than fully realized systems.
@@
  • Multi-Layer Adaptive Frameworks (MLAMs): modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs. MLAFs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
In the following sections, we will analyze each of these conceptual models in detail, evaluating their respective strengths and challenges in the pursuit of a unified framework.
@@ V-C1 Bipedal Foundation Models
Inspired by Robot Foundation Models (RFMs) [96, 115], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion. A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments. Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots. Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [116].
The proposed BFM paradigm would operate in two stages. First, IL on diverse datasets would establish the generalisable foundation. Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot. The potential of this approach is highlighted by recent works, with frameworks like FLaRe [117] enhancing generalisation for long-horizon tasks, MOTO [118] enabling effective offline-to-online adaptation from images, and AdA [119] demonstrating in-context adaptation to novel environments. Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
However, realizing the BFM concept for bipeds presents significant challenges. The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [117, 119]. Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
@@ V-C2 Multi-layered Adaptive Models
As a complementary path to BFMs, we propose the concept of MLAMs. Rather than relying on large-scale pre-training, this conceptual framework would adopt a modular, hierarchical approach. The idea is to extend conventional hierarchical frameworks (discussed in Section III) with explicitly adaptive layers, allowing for the dynamic composition of specialized policies. The core principle of this concept would be modularity, enabling each layer to be independently optimized or replaced and providing interpretable outputs at each stage.
A key feature we envision for MLAMs is their capacity to dynamically compose adaptive modules for each control tier. Each layer processes context-specific inputs and outputs interpretable commands. The HL reasoning layer leverages large pre-trained model such as LLMs and VLMs [120, 121] to parse commands into sub-tasks. For instance, Vision-Language Model Predictive Control [93] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
The mid-level planner selects or synthesises motions via learned motion libraries [58, 44] or DRL-based planners [122]. The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands. These include locomotion primitives like walking and climbing [123], adaptive tracking controllers for whole-body tracking [59], and imitation-based skills such as kicking and dancing [26], by utilizing RL, IL, or model-based methods. This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [123].
However, realizing the MLAM concept would introduce challenges distinct from BFMs. Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [109]. Additionally, the computational latency incurred by HL reasoning modules like LLMs [123] would pose limitations for tasks needing rapid reactions.
@@ VI Conclusion
Despite significant progress in DRL for robotics, a significant gap remains between current achievements and the development of a unified framework that can efficiently handle a wide range of complex tasks. DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have demonstrated success in handling diverse locomotion skills [35], climbing stairs [49], and navigating challenging terrains such as stepping stones [33]. Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
Each framework brings unique strengths to the unified framework. End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility needed for robots to acquire diverse locomotion skills. In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies.
While DRL has enabled remarkable progress, our survey concludes that current frameworks confront key limitations, including the tension between multi-skill generalisation and task-specific precision, the persistent sim-to-real gap, and critical safety concerns. To address these challenges, this survey synthesizes specific pathways for future research and identifies key opportunities for cross-pollination from related fields, such as FMs, loco-manipulation, and quadrupedal robotics. These insights culminate in our proposal of two conceptual frameworks: the BFMs, extending the end-to-end paradigm, and the MLAMs, evolving from the hierarchical approach, which together offer distinct blueprints for the next generation of generalist bipedal controllers.

### v6

@@ V-A7 Constrained learning
While the previously discussed pathways focus on enhancing a robot’s capabilities, a parallel and equally critical frontier is ensuring that these capabilities are exercised safely and reliably. To formally integrate safety, modern approaches can be grouped by how they handle constraints: soft constraints that guide the policy through costs and hard constraints that strictly limit actions [2020_Javier_safereinforcementlearning, 2022_Li_UCB_safteystability_model-based_DRL_PPO].
Soft constraints encourage desirable behaviour and penalise undesirable behaviour without forbidding it. They are well suited to preferences or efficiency goals, for example minimising energy use, limiting peak torques, or promoting smooth motion [2020_Javier_safereinforcementlearning]. Hard constraints are inviolable rules that prevent catastrophic failures. They are essential for enforcing physical limits and protecting the robot and its environment, for example footstep feasibility, contact timing, joint and torque bounds, and collision avoidance [2022_sim2real_footstepconstraint_OSUDRL_specifytouchdownlocation_actorcritic_PPO_LSTM_plannar_transisionmodel_model-predictiveplanning_CNN_predictnexttdlocation, 2021_Casillo_cyberbotics_hierarchyscheme_bezierpolynominals]. A practical way to enforce hard constraints is to use safety filters or shields grounded in control theory, such as control barrier functions and related template model checks [2016_Quan_gaitlibrary].
@@ V-B1 Leveraging foundation models for locomotion learning
The recent rise of Foundation Models (FMs), such as Large Language Models (LLMs) and Vision–Language Models (VLMs), presents a transformative opportunity for bipedal locomotion. Their powerful reasoning capabilities are unlocking new approaches that go beyond traditional control methods, primarily by enabling sophisticated HL task planning and by providing novel solutions to shape the learning process itself, particularly in automated reward design.
As HL planners, FMs provide a reasoning engine that can bridge the gap between abstract human goals and LL motor execution. They can interpret complex linguistic commands or visual scenes and decompose them into a sequence of simpler, actionable commands for an LL policy to follow. This has been demonstrated effectively in legged robotics, where VLMs process raw sensory data to pass structured commands to motor controllers [2024_VLM_PC], creating a seamless link between strategic planning and physical action.
Furthermore, FMs create a significant opportunity to overcome one of the most persistent bottlenecks in DRL: reward design. Instead of tedious manual tuning, LLMs can dynamically generate or refine reward functions based on linguistic descriptions of task success. Research has shown that LLMs can translate human feedback into reward adjustments [2023_kumar_wordscontrol_LLM] or even autonomously adjust rewards and control strategies to self-optimise for diverse terrains [2024_LLM_AnyBipe], drastically reducing human intervention.
The foremost opportunity lies in the deeper synergy between these roles. The integration of the HL symbolic reasoning of FMs with the LL, real-time control of DRL could create a new class of highly adaptive and flexible robots. As this rapidly evolving field progresses, as reviewed in [2023_Fms_survey_robotics_Roya_applicationsChallengeandFuture], we may see a paradigm shift towards more autonomous, self-learning humanoid robots that can understand, reason about, and adapt to the world with minimal human intervention.
@@ V-B2 Loco-manipulation tasks
While achieving stable locomotion is a foundational challenge, a bipedal robot with only a lower body has limited practical utility, as it cannot physically interact with its environment. The evolution of modern humanoids to include complex upper bodies is a critical advancement that has unlocked the opportunity for loco-manipulation—the dynamic integration of movement and object interaction. Achieving such full-body coordination is now a key benchmark for creating truly adaptable systems, with tasks ranging from climbing and using tools to carrying objects while navigating, as highlighted by initiatives like the DARPA Robotics Challenge [2015_DAPRA_Challenge_1].
However, realising this opportunity is a significant challenge. Early studies, such as a ‘box transportation’ framework [2023_dao_OSUDRL_loco-manipulation_hybrid_simtoreal], often rely on inefficient, multi-policy solutions that lack visual perception. Furthermore, dynamically interacting with mobile objects like scooters or balls introduces even greater complexity [2023_baltes_control_scooter, 2023_deepmind_haarnoja_playsoccer].
These difficulties create significant research opportunities. One such opportunity lies in exploring hierarchical control approaches. By decomposing tasks into multiple layers, this method allows for precise, modular control over different components, which can enhance stability and adaptability to environmental variations [2023_template_taskspace_hierarchyscheme_reducedorderstateALIP_learnedhigherlevel_lowlevelinversedynamiccontroller].
Alternatively, a further research opportunity is the development of end-to-end learning frameworks, which offer a more scalable solution. Using techniques like curriculum learning and imitation from human motion-capture data [2021_Diego_deepwalk_curriculum_learning_sim2real, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining, 2023_seo_deep_loco_manipulation, 2024_Zhang_wholebody_adversarial_motion_priors, 2024_speratebody_expressive_allmotion], a single, unified policy can be trained to handle diverse loco-manipulation tasks, representing a promising avenue of research for creating truly versatile agents.
@@ V-B3 Insights from quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end, model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [quadModelBased1, quadModelBased2], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [quadDeformable]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [quadModular, humphreys2024bio], and sophisticated acrobatic motions are achieved through IL [quadImitation2].
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep-planning module is paired with model-based control [quadRLPlanning] within a deep-planning hybrid scheme, and another that combines model-based planning with LL DRL control [quadRLLL, quadmpcref] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs an MPC to generate reference motions, which are then followed by an LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots module [TAMOLS] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Quadruped DRL frameworks are predominantly designed to navigate complex terrains, but efforts to extend their capabilities to other tasks are under way. These include mimicking real animals through motion-capture data and IL [quadAnimal1, quadAnimal2], as well as augmenting quadrupeds with manipulation abilities. This is achieved either by adding a manipulator [quadManip1, quadManip2] or by using the robots’ legs [quadManip3]. Notably, the research presented in [quadManip2] demonstrates that loco-manipulation tasks can be effectively managed using a single, unified, end-to-end framework.
Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers, and high mechanical complexity can be attributed to this disparity between quadruped and bipedal robots. Establishing a unified framework could bridge this gap—an essential step, given the integral role of bipedal robots with upper bodies in developing fully functional humanoid systems. Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
@@ V-C Conceptual models for unified frameworks
Motivated by our survey and the current state of the art, we propose two conceptual models, intended as reference designs, towards a unified locomotion framework. They build on end-to-end and hierarchical paradigms and offer complementary routes to scalable, generalisable architectures, rather than fully realised systems.
@@
  • Multi-Layer Adaptive Models (MLAMs): modular, hierarchical architectures that span from HL planning to LL control, with each layer producing interpretable intermediate outputs. MLAMs are designed to integrate, substitute, and coordinate diverse policies, enabling flexible and adaptive responses across tasks and embodiments.
In the following sections, we will analyse each of these conceptual models in detail, evaluating their respective strengths and challenges in the pursuit of a unified framework.
@@ V-C1 Bipedal foundation models
Inspired by Robot Foundation Models (RFMs) [2023_Fms_survey_robotics_Roya_applicationsChallengeandFuture, 2023_generalpurposerobot_FMs], we propose the concept of BFMs as large-scale, general-purpose models tailored for bipedal locomotion. A BFM would be a large-scale model pre-trained specifically to learn the shared motion priors of dynamic balance and movement across a vast range of bipedal tasks and physical embodiments. Unlike traditional policies trained from scratch, a BFM would provide a foundational understanding of stable locomotion, directly tackling the core difficulties that make bipeds distinct from other robots. Architecturally, we envision such a model comprising a multi-modal embedding module, a shared backbone like a transformer, and an action decoder, drawing inspiration from models like RT-2 [2023_VLA_Rt-2].
The proposed BFM paradigm would operate in two stages. First, IL on diverse datasets would establish the generalisable foundation. Second, DRL would be repurposed as an efficient fine-tuning mechanism to adapt these general priors to the specific, and often unforgiving, dynamics of a physical robot. The potential of this approach is highlighted by recent works, with frameworks like FLaRe [2024_FLaRe_RLfinetune_foundationmodel] enhancing generalisation for long-horizon tasks, MOTO [2024_MOTO_RLfinetune_foundationmodel] enabling effective offline-to-online adaptation from images, and AdA [2023_RL_finetune_RMFs] demonstrating in-context adaptation to novel environments. Collectively, these approaches underscore DRL not only as a simple tuning tool but as a central mechanism for grounding abstract foundation model priors into executable, platform-specific control policies.
However, realising the BFM concept for bipeds presents significant challenges. The DRL fine-tuning stage can be costly and risky on physical hardware, and policies may overfit to narrow dynamics or catastrophically forget the generalisable priors acquired during pre-training [2024_FLaRe_RLfinetune_foundationmodel, 2023_RL_finetune_RMFs]. Furthermore, as detailed in Sections IV-A and V-A4, the scarcity of high-quality, large-scale data remains a fundamental bottleneck, as most existing datasets are human-centric and require significant adaptation before they can be used.
@@ V-C2 Multi-layered adaptive models
As a complementary path to BFMs, we propose the concept of MLAMs. Rather than relying on large-scale pre-training, this conceptual framework would adopt a modular, hierarchical approach. The idea is to extend conventional hierarchical frameworks (discussed in Section III) with explicitly adaptive layers, allowing for the dynamic composition of specialised policies. The core principle of this concept would be modularity, enabling each layer to be independently optimised or replaced and providing interpretable outputs at each stage.
A key feature we envision for MLAMs is their capacity to dynamically compose adaptive modules for each control tier. Each layer processes context-specific inputs and outputs interpretable commands. The HL reasoning layer leverages large pre-trained models such as LLMs and VLMs [2022_DoAsICan, 2022_codeaspolicy_google] to parse commands into sub-tasks. For instance, Vision–Language Model Predictive Control [2024_VLM_PC] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution.
The mid-level planner selects or synthesises motions via learned motion libraries [2021_Kevin_gaitlibrary_heirarchy_reducedorder_gaitlibrary_sim2real, 2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter] or DRL-based planners [2021_Mohammadreza_DRL_improveparameters_forplanner]. The LL control layer comprises various modular controllers, dynamically selected and composed based on task-specific demands. These include locomotion primitives like walking and climbing [2024_yangOu_LLM_hierarchical_quadrupedrobot], adaptive tracking controllers for whole-body tracking [2024_stanford_humanplus_hierarchical_humanoidshadowingimitation], and imitation-based skills such as kicking and dancing [2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task], by utilising RL, IL, or model-based methods. This layered architecture is exemplified by recent work on quadrupedal robots, where LLMs are used to translate HL commands into robust and flexible real-world behaviours [2024_yangOu_LLM_hierarchical_quadrupedrobot].
However, realising the MLAM concept would introduce challenges distinct from BFMs. Such a framework would depend heavily on real-time multi-modal perception, which complicates data alignment across layers with differing timescales and abstraction levels [TAMOLS]. Additionally, the computational latency incurred by HL reasoning modules like LLMs [2024_yangOu_LLM_hierarchical_quadrupedrobot] would pose limitations for tasks needing rapid reactions.
@@ VI Conclusion
Despite significant progress in DRL for robotics, a substantial gap remains between current achievements and the development of a unified framework capable of efficiently handling a wide range of complex tasks. DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have demonstrated success in handling diverse locomotion skills [2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining], climbing stairs [2021_siekmann_blind_DRL_stair], and navigating challenging terrains such as stepping stones [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark]. Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
Each framework contributes unique strengths to the pursuit of a unified framework. End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility required for robots to acquire diverse locomotion skills. In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies.
While DRL has enabled remarkable progress, our survey concludes that current frameworks face key limitations, including the tension between multi-skill generalisation and task-specific precision, the persistent sim-to-real gap, and critical safety concerns. To address these challenges, this survey synthesises specific pathways for future research and identifies key opportunities for cross-pollination from related fields, such as FMs, loco-manipulation, and quadrupedal robotics. These insights culminate in our proposal of two conceptual frameworks: the BFMs, extending the end-to-end paradigm, and the MLAMs, evolving from the hierarchical approach, which together offer distinct blueprints for the next generation of generalist bipedal controllers.